The normal duration of physics practical period in Narayana colleges 100 minutes Express this period in micro centuries
 

Grade:8

1 Answers

Eshan
5 years ago
Dear student,

1 micro centuary=10^{-6}century
1 century=100 years=
1 year=365 days
1 day=24 hours
1 hour=60 minutes

Therefore 1 mirco century=10^{-6}\times 100\times 365\times 24\times 60minute
Therefore 1 minute=\dfrac{1}{10^{-6}\times 100\times 365\times 24\times 60}microcenturies

100 minutes=\dfrac{100}{10^{-6}\times 100\times 365\times 24\times 60}microcenturies=1.9 microcenturies